Incompressible fluid flow field2. (a) An incompressible fluid flow field is given as Vx = x2+y+z2 and Vy=xy+yz+z, what is V?=? that satisfies continuity equ
Answer #1

oz for satisfying continuity equation. on the org + SME Co. - Von = x2 +42+72 av = 25e toto Vy = xy tyztz, avy _ xtz to Dyst2a=24 al y = lsx = 2 by 24 =2 y.dg - 4 = 742 + f(x) - -- Vy = - 4 > oq = -4nda 393 - 4 m2 + f(y)+c 4 = -2*2 + f (y) + C. - Now